OpenStudy (anonymous):

\[Solve ~the~equation ~~~tanx=3cotx, giving~all~solutions ~btween~0 ~and~36\]

OpenStudy (anonymous):

on simplification we get cos2x=cos120 thus x=n*pi +60,n*pi - 60 here n is an integrer

OpenStudy (anonymous):

\[\tan^2x=3\]

OpenStudy (anonymous):

tan^2x = 3 (tanx - sqrt3)(tanx + sqrt3) = 0

OpenStudy (anonymous):

tanx= sqrt(3) and tanx =-sqrt(3) \[x=n \pi \pm \]

OpenStudy (anonymous):

I got tan^2x=3... But I'm not sure how to get the values between 0 and 36.

OpenStudy (anonymous):

360*

OpenStudy (anonymous):

\[x=n \pi \pm \pi/3\]

OpenStudy (anonymous):

Use the Unit Circle #TheAmericanWay

OpenStudy (anonymous):

Or, follow NSB heh

OpenStudy (anonymous):

ans are \[\pi/3 ,2\pi/3 ,4\pi/3, 5 \pi/3\]

OpenStudy (anonymous):

Can't quite follow. Can you show me an example of how to get the answers... They're in degrees I think...